Find the algebraic expresson for the quantity in italics using the given information:
The rate, given that the distance is 200 feet and the time is x + 3 seconds. (rate is italicized)

Answer by richwmiller(17219)

We have the distance and the time.
What is the formula that connect rate time and distance?
r*t=d
r=d/t
r=200/(x+3)
The result will be feet per second.